# VideoBadminton: A Video Dataset for Badminton Action Recognition

### v1

@@ B.0.1. R(2+1)D
type
ResNet2Plus1d
depth
34
pretrained
None
pretrained2d
False
norm_eval
False
conv_cfg
{ Conv2plus1d}
norm_cfg
{ SyncBN, requires_grad: True, eps: 0.001}
conv1_kernel
(3, 7, 7)
conv1_stride_t
1
pool1_stride_t
1
inflate
(1, 1, 1, 1)
spatial_strides
(1, 2, 2, 2)
temporal_strides
(1, 2, 2, 2)
zero_init_residual
False
@@
type
I3DHead
num_classes
18
in_channels
512
spatial_type
avg
dropout_ratio
0.5
init_std
0.01
average_clips
prob
@@
type
ActionDataPreprocessor
mean
[123.675, 116.28, 103.53]
std
[58.395, 57.12, 57.375]
format_shape
NCTHW
@@
1
DecordInit, io_backend: disk
2
SampleFrames, clip_len: 8, frame_interval: 8, num_clips: 1
3
DecordDecode
4
Resize, scale: (-1, 256)
5
RandomResizedCrop
6
Resize, scale: (224, 224), keep_ratio: False
7
Flip, flip_ratio: 0.5
8
FormatShape, input_format: NCTHW
9
PackActionInputs
@@ B.0.2. SlowFast
type
ResNet3dSlowFast
pretrained
None
resample_rate
8
speed_ratio
8
channel_ratio
8
@@
type
resnet3d
depth
50
pretrained
None
lateral
True
conv1_kernel
(1, 7, 7)
dilations
(1, 1, 1, 1)
conv1_stride_t
1
pool1_stride_t
1
inflate
(0, 0, 1, 1)
norm_eval
False
@@
type
resnet3d
depth
50
pretrained
None
lateral
False
base_channels
8
conv1_kernel
(5, 7, 7)
conv1_stride_t
1
pool1_stride_t
1
norm_eval
False
@@
type
SlowFastHead
in_channels
2304
num_classes
18
spatial_type
avg
dropout_ratio
0.5
average_clips
prob
@@
type
ActionDataPreprocessor
mean
[123.675, 116.28, 103.53]
std
[58.395, 57.12, 57.375]
format_shape
NCTHW
@@
1
DecordInit, io_backend: disk
2
SampleFrames, clip_len: 32, frame_interval: 2, num_clips: 1
3
DecordDecode
4
Resize, scale: (-1, 256)
5
RandomResizedCrop
6
Resize, scale: (224, 224), keep_ratio: False
7
Flip, flip_ratio: 0.5
8
FormatShape, input_format: NCTHW
9
PackActionInputs
@@ B.0.3. TimeSformer
type
TimeSformer
num_frames
8
img_size
224
patch_size
16
embed_dims
768
in_channels
3
dropout_ratio
0.0
transformer_layers
None
attention_type
divided_space_time
norm_cfg
{ LN, eps: 1e-06}
@@
type
TimeSformerHead
num_classes
18
in_channels
768
average_clips
prob
@@
type
ActionDataPreprocessor
mean
[127.5, 127.5, 127.5]
std
[127.5, 127.5, 127.5]
format_shape
NCTHW
@@
1
DecordInit, io_backend: disk
2
SampleFrames, clip_len: 8, frame_interval: 2, num_clips: 1
3
DecordDecode
4
RandomRescale, scale_range: (256, 320)
5
RandomCrop, size: 224
6
Flip, flip_ratio: 0.5
7
FormatShape, input_format: NCTHW
8
PackActionInputs
@@ B.0.4. PoseC3D
type
ResNet3dSlowOnly
depth
50
pretrained
None
in_channels
17
base_channels
32
num_stages
3
out_indices
(2, )
stage_blocks
(4, 6, 3)
conv1_stride_s
1
pool1_stride_s
1
inflate
(0, 1, 1)
spatial_strides
(2, 2, 2)
temporal_strides
(1, 1, 2)
dilations
(1, 1, 1)
@@
type
I3DHead
in_channels
512
num_classes
18
dropout_ratio
0.5
average_clips
prob
@@
1
UniformSampleFrames, clip_len: 48
2
PoseDecode
3
PoseCompact, hw_ratio: 1.0, allow_imgpad: True
4
Resize, scale: (-1, 64)
5
RandomResizedCrop, area_range: (0.56, 1.0)
6
Resize, scale: (56, 56), keep_ratio: False
@@
9
FormatShape, input_format: NCTHW_Heatmap
10
PackActionInputs
@@ B.0.5. Swim
type
SwinTransformer3D
arch
small
pretrained2d
True
patch_size
(2, 4, 4)
window_size
(8, 7, 7)
mlp_ratio
4.0
qkv_bias
True
qk_scale
None
drop_rate
0.0
attn_drop_rate
0.0
drop_path_rate
0.2
patch_norm
True
@@
type
ActionDataPreprocessor
mean
[123.675, 116.28, 103.53]
std
[58.395, 57.12, 57.375]
format_shape
NCTHW
@@
type
I3DHead
in_channels
768
num_classes
18
spatial_type
avg
dropout_ratio
0.5
average_clips
prob
@@
1
type: DecordInit, io_backend: disk
2
type: SampleFrames, clip_len: 32, frame_interval: 2, num_clips: 1
3
type: DecordDecode
4
type: Resize, scale: (-1, 256)
5
type: RandomResizedCrop
6
type: Resize, scale: (224, 224), keep_ratio: False
7
type: Flip, flip_ratio: 0.5
8
type: FormatShape, input_format: NCTHW
9
type: PackActionInputs
@@ B.0.6. Mvit-V2
type
MViT
arch
small
drop_path_rate
0.2
@@
type
ActionDataPreprocessor
mean
[114.75, 114.75, 114.75]
std
[57.375, 57.375, 57.375]
format_shape
NCTHW
@@
  - type
MixupBlending, alpha: 0.8, num_classes: 18
  - type
CutmixBlending, alpha: 1, num_classes: 18
@@
type
MViTHead
in_channels
768
num_classes
18
label_smooth_eps
0.1
average_clips
prob
@@
1
DecordInit, io_backend: disk
2
SampleFrames, clip_len: 16, frame_interval: 4, num_clips: 1
3
DecordDecode
4
Resize, scale: (-1, 256)
5
PytorchVideoWrapper, op: RandAugment, magnitude: 7, num_layers: 4
6
RandomResizedCrop
7
Resize, scale: (224, 224), keep_ratio: False
8
Flip, flip_ratio: 0.5
9
RandomErasing, erase_prob: 0.25, mode: ’rand’
10
FormatShape, input_format: NCTHW
11
PackActionInputs
@@ B.0.7. ST-GCN
type
STGCN
graph_cfg
{layout: ’coco’, mode: ’stgcn_spatial’}
@@
type
GCNHead
num_classes
18
in_channels
256
@@
1
PreNormalize2D
2
GenSkeFeat, dataset: ’coco’, feats: [’b’]
3
UniformSampleFrames, clip_len: 100
4
PoseDecode
5
FormatGCNInput, num_person: 2
6
PackActionInputs

### v2

# Benchmarking Badminton Action Recognition with a New Fine-Grained Dataset

@@ B.0.1. R(2+1)D
type ResNet2Plus1d
depth 34
pretrained None
pretrained2d False
norm_eval False
conv_cfg { Conv2plus1d}
norm_cfg { SyncBN, requires_grad: True, eps: 0.001}
conv1_kernel (3, 7, 7)
conv1_stride_t 1
pool1_stride_t 1
inflate (1, 1, 1, 1)
spatial_strides (1, 2, 2, 2)
temporal_strides (1, 2, 2, 2)
zero_init_residual False
@@
type I3DHead
num_classes 18
in_channels 512
spatial_type avg
dropout_ratio 0.5
init_std 0.01
average_clips prob
@@
type ActionDataPreprocessor
mean [123.675, 116.28, 103.53]
std [58.395, 57.12, 57.375]
format_shape NCTHW
@@
1 DecordInit, io_backend: disk
2 SampleFrames, clip_len: 8, frame_interval: 8, num_clips: 1
3 DecordDecode
4 Resize, scale: (-1, 256)
5 RandomResizedCrop
6 Resize, scale: (224, 224), keep_ratio: False
7 Flip, flip_ratio: 0.5
8 FormatShape, input_format: NCTHW
9 PackActionInputs
@@ B.0.2. SlowFast
type ResNet3dSlowFast
pretrained None
resample_rate 8
speed_ratio 8
channel_ratio 8
@@
type resnet3d
depth 50
pretrained None
lateral True
conv1_kernel (1, 7, 7)
dilations (1, 1, 1, 1)
conv1_stride_t 1
pool1_stride_t 1
inflate (0, 0, 1, 1)
norm_eval False
@@
type resnet3d
depth 50
pretrained None
lateral False
base_channels 8
conv1_kernel (5, 7, 7)
conv1_stride_t 1
pool1_stride_t 1
norm_eval False
@@
type SlowFastHead
in_channels 2304
num_classes 18
spatial_type avg
dropout_ratio 0.5
average_clips prob
@@
type ActionDataPreprocessor
mean [123.675, 116.28, 103.53]
std [58.395, 57.12, 57.375]
format_shape NCTHW
@@
1 DecordInit, io_backend: disk
2 SampleFrames, clip_len: 32, frame_interval: 2, num_clips: 1
3 DecordDecode
4 Resize, scale: (-1, 256)
5 RandomResizedCrop
6 Resize, scale: (224, 224), keep_ratio: False
7 Flip, flip_ratio: 0.5
8 FormatShape, input_format: NCTHW
9 PackActionInputs
@@ B.0.3. TimeSformer
type TimeSformer
num_frames 8
img_size 224
patch_size 16
embed_dims 768
in_channels 3
dropout_ratio 0.0
transformer_layers None
attention_type divided_space_time
norm_cfg { LN, eps: 1e-06}
@@
type TimeSformerHead
num_classes 18
in_channels 768
average_clips prob
@@
type ActionDataPreprocessor
mean [127.5, 127.5, 127.5]
std [127.5, 127.5, 127.5]
format_shape NCTHW
@@
1 DecordInit, io_backend: disk
2 SampleFrames, clip_len: 8, frame_interval: 2, num_clips: 1
3 DecordDecode
4 RandomRescale, scale_range: (256, 320)
5 RandomCrop, size: 224
6 Flip, flip_ratio: 0.5
7 FormatShape, input_format: NCTHW
8 PackActionInputs
@@ B.0.4. PoseC3D
type ResNet3dSlowOnly
depth 50
pretrained None
in_channels 17
base_channels 32
num_stages 3
out_indices (2, )
stage_blocks (4, 6, 3)
conv1_stride_s 1
pool1_stride_s 1
inflate (0, 1, 1)
spatial_strides (2, 2, 2)
temporal_strides (1, 1, 2)
dilations (1, 1, 1)
@@
type I3DHead
in_channels 512
num_classes 18
dropout_ratio 0.5
average_clips prob
@@
1 UniformSampleFrames, clip_len: 48
2 PoseDecode
3 PoseCompact, hw_ratio: 1.0, allow_imgpad: True
4 Resize, scale: (-1, 64)
5 RandomResizedCrop, area_range: (0.56, 1.0)
6 Resize, scale: (56, 56), keep_ratio: False
@@
9 FormatShape, input_format: NCTHW_Heatmap
10 PackActionInputs
@@ B.0.5. Swim
type SwinTransformer3D
arch small
pretrained2d True
patch_size (2, 4, 4)
window_size (8, 7, 7)
mlp_ratio 4.0
qkv_bias True
qk_scale None
drop_rate 0.0
attn_drop_rate 0.0
drop_path_rate 0.2
patch_norm True
@@
type ActionDataPreprocessor
mean [123.675, 116.28, 103.53]
std [58.395, 57.12, 57.375]
format_shape NCTHW
@@
type I3DHead
in_channels 768
num_classes 18
spatial_type avg
dropout_ratio 0.5
average_clips prob
@@
1 type: DecordInit, io_backend: disk
2 type: SampleFrames, clip_len: 32, frame_interval: 2, num_clips: 1
3 type: DecordDecode
4 type: Resize, scale: (-1, 256)
5 type: RandomResizedCrop
6 type: Resize, scale: (224, 224), keep_ratio: False
7 type: Flip, flip_ratio: 0.5
8 type: FormatShape, input_format: NCTHW
9 type: PackActionInputs
@@ B.0.6. Mvit-V2
type MViT
arch small
drop_path_rate 0.2
@@
type ActionDataPreprocessor
mean [114.75, 114.75, 114.75]
std [57.375, 57.375, 57.375]
format_shape NCTHW
@@
    - type MixupBlending, alpha: 0.8, num_classes: 18
    - type CutmixBlending, alpha: 1, num_classes: 18
@@
type MViTHead
in_channels 768
num_classes 18
label_smooth_eps 0.1
average_clips prob
@@
1 DecordInit, io_backend: disk
2 SampleFrames, clip_len: 16, frame_interval: 4, num_clips: 1
3 DecordDecode
4 Resize, scale: (-1, 256)
5 PytorchVideoWrapper, op: RandAugment, magnitude: 7, num_layers: 4
6 RandomResizedCrop
7 Resize, scale: (224, 224), keep_ratio: False
8 Flip, flip_ratio: 0.5
9 RandomErasing, erase_prob: 0.25, mode: ’rand’
10 FormatShape, input_format: NCTHW
11 PackActionInputs
@@ B.0.7. ST-GCN
type STGCN
graph_cfg {layout: ’coco’, mode: ’stgcn_spatial’}
@@
type GCNHead
num_classes 18
in_channels 256
@@
1 PreNormalize2D
2 GenSkeFeat, dataset: ’coco’, feats: [’b’]
3 UniformSampleFrames, clip_len: 100
4 PoseDecode
5 FormatGCNInput, num_person: 2
6 PackActionInputs